OpenStudy (anonymous):

The ratio of copper and iron in an alloy is 4:5 .If there is 25% impurity in copper then find the amount of impurity in 18 kg of the alloy.

OpenStudy (shane_b):

There are 4+5=9 parts that make up the alloy. The copper portion is 0.75*4=3 parts and the iron is 5 parts. Therefore, 1 part in 9 consists of impurities. \[\frac{1}{9} * 18 = 2kg\]